The following are selected 2020 transactions of Indigo Corporation.

Sept. 1 Purchased inventory from Encino Company on account for $55,600. Indigo records purchases gross and uses a periodic inventory system.
Oct. 1 Issued a $55,600, 12-month, 8% note to Encino in payment of account.
Oct. 1

Borrowed $55,600 from the Shore Bank by signing a 12-month, zero-interest-bearing $59,960 note.

Compute the total net liability to be reported on the December 31 balance sheet for:

(1)

The interest-bearing note

$enter a dollar amount
(2)

The zero-interest-bearing note

$enter a dollar amount
Answer #1

1. Total net liability (the interest bearing note) = 55600 + 55600*8%*3/12 = $56712

2. Total net liability (the zero interest bearing note) = 55600 + (59960-55600)*3/12 = $56690